An express train passes through a station that is 204 m long. It enters with an initial speed of 24 m/s and decelerates at a rate of 0.11 m/s2 as it goes through.

a. During what time interval, is the front of the train within the station?

b. How fast is the train moving when its front leaves the station?

c. If the train is 134 m long, how long after the front of the train enters the station does the end of the train leave the station?

d. What is the velocity of the end of the train as it leaves?
Answer #1

a)Using kinematics equation s=vj+za 204 = 241 +3(-0.11) 0.055t - 24t + 204 = 0 By solving the equation t = 8.67s t=427.75 Aga